OpenStudy (anonymous):

Given the equation x^2+y^2+4x+16y-15=0, (a) find the center and radius; (b) find the intercepts, if any. Please show all of your work.

OpenStudy (anonymous):

(x+2)^2 - 4 + (y+8)^2 - 64 = 15 (x+2)^2 + (y+8)^2 = 83 (a) centre = (-2,-8) and radius = sqrt(83) sorry gotta go - cant complete (

OpenStudy (anonymous):